OpenStudy (anonymous):

how would you solve 6p+9=5p

OpenStudy (anonymous):



OpenStudy (anonymous):

We already answered this exact question. @RadEn and myself?

OpenStudy (anonymous):

can you explain how you got the -9

OpenStudy (anonymous):



OpenStudy (anonymous):

I'm sorry I am not understanding because of the 6-5 why is that not in the answer

OpenStudy (anonymous):

http://openstudy.com/users/corrina.howerton#/updates/5158a726e4b07077e0c09c7c @RadEn explained like this in previous post. first, subtract by 5p to both sides 6p+9-5p = 5p - 5p simplify, will be : p+9 = 0 then subtract both sides by -9 p+9-9 = 0-9 p = -9

OpenStudy (anonymous):

there is no 6 - 5? I said 6p - 5p = 1p = p

OpenStudy (anonymous):

that was the 6 I was talking about, I think I understand now because now the 1p is = to just p with out the number in front

OpenStudy (anonymous):

yes precisely. The goal is to isolate the variable like how i did in this example. i ISOLATED p.

OpenStudy (anonymous):

look over the work i did a few times if that might help :D

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

No problem :D